# Quantum control of continuous systems via nonharmonic potential modulation

### v3

@@ 4.3 Coherent couplings
Coherent coupling to other degrees of freedom might cause an additional experimental challenge. The exemplary case of such a noisy channel occurs in one of the setups we consider and is due to the dimensionality of the realistic optical potential. The three-dimensional tweezer potential is necessarily nonseparable and, depending on the axis of control, axial or radial, a one-dimensional approximation may be less strictly satisfied [121]. Restricting control to only one direction unavoidably leads to excitation in other ones. However, the fidelity loss due to these excitations can be alleviated through additional optimization in fully three-dimensional potential and additional optimized position and depth controls. The additional control then effectively deexcites the coupled degrees of freedom. This optimization can be done via a staged procedure, where the initial guess for a higher-dimensional system comes from a reduced-dimensional control optimization. Such an optimization procedure can be utilized in setups in which coherent dynamics beyond one-dimensional approximation can be solved.
@@
Figure 10: Fock $\psi_{1}(x)$ state preparation protocol with the optimally controlled Gaussian potential characterized with $\eta=0.25$ and $\lambda=0.167$ for (a) one-, (b) two-, and (c) three-dimensional geometries. Achieved fidelities $\mathcal{F}$ read (a) $99.7\%$, (b) $99.4\%$, and (c) $99.0\%$.
@@
and optimized through a staged procedure: the optimized control in the one-dimensional problem is a starting guess for the two-dimensional control; analogously for the three-dimensional case. In Fig. 10 we plot the results of such an optimization with the snapshots of Wigner functions associated with reduced density matrices,
@@ 5 Conclusions & outlook
The Fourier transforms of all presented controls are presented in Appendix D, while the depth modulations not shown in the main text are plotted in Appnedix E. Data analysis and simulation codes are available on Zenodo upon reasonable request [125].
We thank R. G. Cortiñas, N. E. Frattini, S. Muleady, A. M. Rey, and P. Zoller for helpful discussions. The computational results presented here have been achieved in part using the LEO HPC infrastructure of the University of Innsbruck. P.T.G. and O.R.-I. have been supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 863132 (IQLev) and by the European Research Council (ERC) under the grant agreement No. 951234 (Q-Xtreme ERC-2020-SyG). P.T.G. was partially supported by the project CZ.02.01.01/00/22_008/0004649 (QUEENTEC) of EU. H.P. has been supported through an ERC Starting grant QARA (grant no. 101041435) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101079862 (PASQuanS2). C.A.R. has been supported by NSF PHY-2317149, NSF QLCI award OMA 2016244, the US Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the Baur-SPIE Chair at JILA.
@@ Appendix A Single-well two-level unitary implementations
Figure A1: Evolution of selected orthogonal states in a cosine potential characterized by $\eta=0.2$ and with optimally controlled displacement $u(\tau)$ and intensity $a(\tau)$ that are constrained through the flux control of a flux-tunable transmon with $d=0.8$. The solid line shows $u(\tau)$ and the protocol lasts $\tau_{\text{max}}/2\pi=9$. (a-b) Hadamard unitary within a subspace spanned by |$\ket{0}$ and |$\ket{1}$ Fock states. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 99.99\%$. (c-d) $\sigma_{y}$ unitary within a subspace spanned by GKP states, $\psi^{\pm}=\psi_{\text{GKP}}(x\pm d_{3}/4)$ with $r=0.7$, $(d_{1},d_{2},d_{3})=(-\sqrt{6\pi},0,\sqrt{6\pi})$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 89.5\%$. (e-f) $\sigma_{x}$ unitary within a subspace spanned by four-legged-cat states states with $\beta=2$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 94.0\%$.
@@ Appendix B Double-well two-level unitary implementations
In Fig. B1, we present examples of the implementation of selected unitaries within a two-level subspace spanned by Kerr-cat states, $\psi^{\pm}(x)=[\psi_{0}(x+s/2)\pm i\psi_{0}(x-s/2)]/\sqrt{2}$ with $s=9$ via double-well Gaussian potential with $\eta=0.25$.
Figure B1: Evolution of selected orthogonal Kerr-cat states in a double-well Gaussian potential with optimally controlled position displacements $u_{1}(\tau)$ and $u_{2}(\tau)$. The solid lines show $u_{1}(\tau)$ and $u_{2}(\tau)$ and the protocol lasts $\tau_{\text{max}}/2\pi=6$. (a-b) $\sigma_{x}$ unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 99.2\%$. (c-d) $\sigma_{x}$ unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 90.9\%$. (e-f) Hadamard unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 95.9\%$.
@@ Appendix C Selective stealing protocols
Figure C1: Spatial separation of orthogonal states $\psi^{\pm}(x+s/2)\rightarrow\psi^{\pm}(x\mp s/2)$ with $s=9$ for (a-b) $\psi^{+}(x)=\psi_{0}(x)$ and $\psi^{-}(x)=\psi_{1}(x)$, (c-d) $\psi^{\pm}(x)=[\psi_{0}(x)\pm\psi_{1}(x)]/\sqrt{2}$, and (e-f) $\psi^{\pm}(x)=[\psi_{0}(x)\pm i\psi_{1}(x)]/\sqrt{2}$. Fidelities read (a-b) $\mathcal{F}_{\hat{U}}\approx 98.9\%$, (c-d) $\mathcal{F}_{\hat{U}}\approx 98.9\%$, and (e-f) $\mathcal{F}_{\hat{U}}\approx 97.4\%$
@@ Appendix D Fourier transforms of control functions
for all the presented protocols. Specifically, Figs. D-1D-3, and D-7 pertain to the figures from the main text, while Figs. D-2D-4D-A1D-B1D-6D-C1, and D-10 are associated to the Supplemental Materials.
@@
Figure D-3: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 3 in the main text. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
Figure D-4: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 4. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
@@
Figure D-7: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 7 in the main text. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
Figure D-10: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 10. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
@@ Appendix E Depth modulations
Some of the protocols, presented in both the main text and the supplementary materials, utilized additional optimal control of the potential’s depth, $a(\tau)$. In Figs. E-3, E-4, and E-6, these control functions are shown.
Figure E-3: Depth control functions $a(\tau)$ from Fig. 3.
Figure E-4: Depth control functions $a(\tau)$ from Fig. 4.

### v4

@@ 4.3 Coherent couplings
Coherent coupling to other degrees of freedom might cause an additional experimental challenge. The exemplary case of such a noisy channel occurs in one of the setups we consider and is due to the dimensionality of the realistic optical potential. The three-dimensional tweezer potential is necessarily nonseparable and, depending on the axis of control, axial or radial, a one-dimensional approximation may be less strictly satisfied [123]. Restricting control to only one direction unavoidably leads to excitation in other ones. However, the fidelity loss due to these excitations can be alleviated through additional optimization in a fully three-dimensional potential and additional optimized position and depth controls. The additional control then effectively deexcites the coupled degrees of freedom. This optimization can be done via a staged procedure, where the initial guess for a higher-dimensional system comes from a reduced-dimensional control optimization. Such an optimization procedure can be utilized in setups in which coherent dynamics beyond one-dimensional approximation can be solved.
@@
Figure 11: Fock $\psi_{1}(x)$ state preparation protocol with the optimally controlled Gaussian potential characterized with $\eta=0.25$ and $\lambda=0.167$ for (a) one-, (b) two-, and (c) three-dimensional geometries. Achieved fidelities $\mathcal{F}$ read (a) $99.7\%$, (b) $99.4\%$, and (c) $99.0\%$. See the caption of Fig. 1 for subplot and curve legend details.
@@
and optimized through a staged procedure: the optimized control in the one-dimensional problem is a starting guess for the two-dimensional control; analogously for the three-dimensional case. In Fig. 11, we plot the results of such an optimization with the snapshots of Wigner functions associated with reduced density matrices,
@@ 5 Conclusions & outlook
The Fourier transforms of all presented controls are presented in Appendix D, while the depth modulations not shown in the main text are plotted in Appendix E. Data analysis and simulation codes are available on Zenodo [127].
We thank R. G. Cortiñas, N. E. Frattini, S. Muleady, A. M. Rey, and P. Zoller for helpful discussions. The computational results presented here have been achieved in part using the LEO HPC infrastructure of the University of Innsbruck. P.T.G. and O.R.-I. have been supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 863132 (IQLev) and by the European Research Council (ERC) under the grant agreement No. 951234 (Q-Xtreme ERC-2020-SyG). P.T.G. was partially supported by the project CZ.02.01.01/00/22_008/0004649 (QUEENTEC) of EU and project 23-06308S of the Czech Science Foundation. H.P. has been supported through an ERC Starting grant QARA (grant no. 101041435) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101079862 (PASQuanS2). C.A.R. has been supported by NSF PHY-2317149, NSF QLCI award OMA 2016244, the US Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the Baur-SPIE Chair at JILA.
@@ Appendix A Single-well two-level unitary implementations
[…]
@@ Appendix B Double-well two-level unitary implementations
In Fig. B1, we present examples of the implementation of selected unitaries within a two-level subspace spanned by Kerr-cat states, $\psi^{\pm}(x)=[\psi_{0}(x+s/2)\pm i\psi_{0}(x-s/2)]/\sqrt{2}$ with $s=9$ via a double-well Gaussian potential with $\eta=0.25$.
Figure B1: Evolution of selected orthogonal Kerr-cat states in a double-well Gaussian potential with optimally controlled position displacements $u_{1}(\tau)$ and $u_{2}(\tau)$. The solid lines show $u_{1}(\tau)$ and $u_{2}(\tau)$ and the protocol lasts $\tau_{\text{max}}/2\pi=6$. (a-b) $\sigma_{x}$ unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 99.2\%$. (c-d) $\sigma_{x}$ unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 90.9\%$. (e-f) Hadamard unitary. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 95.9\%$. See the caption of Fig. 1 for subplot and curve legend details.
@@ Appendix C Selective stealing protocols
Figure C1: Spatial separation of orthogonal states $\psi^{\pm}(x+s/2)\rightarrow\psi^{\pm}(x\mp s/2)$ with $s=9$ for (a-b) $\psi^{+}(x)=\psi_{0}(x)$ and $\psi^{-}(x)=\psi_{1}(x)$, (c-d) $\psi^{\pm}(x)=[\psi_{0}(x)\pm\psi_{1}(x)]/\sqrt{2}$, and (e-f) $\psi^{\pm}(x)=[\psi_{0}(x)\pm i\psi_{1}(x)]/\sqrt{2}$. Fidelities read (a-b) $\mathcal{F}_{\hat{U}}\approx 98.9\%$, (c-d) $\mathcal{F}_{\hat{U}}\approx 98.9\%$, and (e-f) $\mathcal{F}_{\hat{U}}\approx 97.4\%$. See the caption of Fig. 1 for subplot and curve legend details.
@@ Appendix D Fourier transforms of control functions
for all the presented protocols. Specifically, Figs. D-1D-5, and D-9 pertain to the figures from the main text, while Figs. D-2D-6D-A1D-B1D-8D-C1, and D-11 are associated to the Supplemental Materials.
@@
Figure D-3: Fourier transform of displacement control function $u(\tau)$ from Fig. 3(a). The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
Figure D-5: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 5 in the main text. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
@@
Figure D-8: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 8. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
Figure D-9: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 9 in the main text. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
Figure D-11: Fourier transforms of displacement control functions $u(\tau)$ from Fig. 11. The vertical axis is in arbitrary units, normalized to the highest frequency contribution.
@@ Appendix E Depth modulations
Some of the protocols, presented in both the main text and the supplementary materials, utilized additional optimal control of the potential’s depth, $a(\tau)$. In Figs. E-5, E-6, and E-8, these control functions are shown.
Figure E-5: Depth control functions $a(\tau)$ from Fig. 5.